# Graph Signal Processing for Global Stock Market Volatility Forecasting

### v1

@@ 5.3 Experiment Results
v-GSPHAR
vs GSPHAR
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.056}$ $\mathbf{0.001}$ $\mathbf{6.550}$ $\mathbf{0}$ $\mathbf{4.696}$ $\mathbf{0}$ $\mathbf{2.066}$ $\mathbf{0.019}$ $\mathbf{3.741}$ $\mathbf{0}$
AEX $\mathbf{3.243}$ $\mathbf{0.001}$ $\mathbf{4.406}$ $\mathbf{0}$ $\mathbf{4.004}$ $\mathbf{0}$ $\mathbf{2.302}$ $\mathbf{0.011}$ $\mathbf{3.787}$ $\mathbf{0}$
BFX $\mathbf{2.278}$ $\mathbf{0.011}$ $\mathbf{3.548}$ $\mathbf{0}$ $\mathbf{2.855}$ $\mathbf{0.002}$ $\mathbf{2.890}$ $\mathbf{0.002}$ $0.690$ $0.245$
STOXX50E $\mathbf{3.488}$ $\mathbf{0}$ $\mathbf{6.351}$ $\mathbf{0}$ $\mathbf{4.853}$ $\mathbf{0}$ $\mathbf{2.968}$ $\mathbf{0.002}$ $\mathbf{3.426}$ $\mathbf{0}$
IBEX $\mathbf{2.729}$ $\mathbf{0.003}$ $\mathbf{7.057}$ $\mathbf{0}$ $\mathbf{5.014}$ $\mathbf{0}$ $\mathbf{3.364}$ $\mathbf{0}$ $\mathbf{3.519}$ $\mathbf{0}$
GDAXI $\mathbf{3.603}$ $\mathbf{0}$ $\mathbf{5.551}$ $\mathbf{0}$ $\mathbf{4.983}$ $\mathbf{0}$ $\mathbf{2.795}$ $\mathbf{0.003}$ $\mathbf{3.988}$ $\mathbf{0}$
AORD $\mathbf{2.529}$ $\mathbf{0.006}$ $\mathbf{2.119}$ $\mathbf{0.017}$ $\mathbf{2.913}$ $\mathbf{0.002}$ $\mathbf{2.724}$ $\mathbf{0.003}$ $\mathbf{3.180}$ $\mathbf{0.001}$
FTSE $\mathbf{3.048}$ $\mathbf{0.001}$ $\mathbf{3.396}$ $\mathbf{0}$ $\mathbf{3.567}$ $\mathbf{0}$ $\mathbf{3.254}$ $\mathbf{0.001}$ $\mathbf{2.718}$ $\mathbf{0.003}$
MXX $0.610$ $0.271$ $\mathbf{3.835}$ $\mathbf{0}$ $1.117$ $0.132$ $\mathbf{7.176}$ $\mathbf{0}$ $1.403$ $0.080$
IXIC $\mathbf{2.009}$ $\mathbf{0.022}$ $\mathbf{2.575}$ $\mathbf{0.005}$ $\mathbf{2.313}$ $\mathbf{0.010}$ $\mathbf{2.496}$ $\mathbf{0.006}$ $\mathbf{1.951}$ $\mathbf{0.026}$
SSMI $\mathbf{3.234}$ $\mathbf{0.001}$ $\mathbf{4.484}$ $\mathbf{0}$ $\mathbf{3.569}$ $\mathbf{0}$ $-0.149$ $0.559$ $\mathbf{3.476}$ $\mathbf{0}$
SPX $\mathbf{3.327}$ $\mathbf{0}$ $\mathbf{3.965}$ $\mathbf{0}$ $\mathbf{3.998}$ $\mathbf{0}$ $1.370$ $0.085$ $\mathbf{3.262}$ $\mathbf{0.001}$
RUT $\mathbf{1.667}$ $\mathbf{0.048}$ $\mathbf{3.054}$ $\mathbf{0.001}$ $\mathbf{2.246}$ $\mathbf{0.012}$ $\mathbf{2.920}$ $\mathbf{0.002}$ $1.387$ $0.083$
DJI $\mathbf{3.438}$ $\mathbf{0}$ $\mathbf{3.570}$ $\mathbf{0}$ $\mathbf{4.055}$ $\mathbf{0}$ $1.586$ $0.056$ $\mathbf{3.026}$ $\mathbf{0.001}$
BSESN $\mathbf{4.805}$ $\mathbf{0}$ $\mathbf{6.721}$ $\mathbf{0}$ $\mathbf{4.738}$ $\mathbf{0}$ $0.762$ $0.223$ $\mathbf{4.901}$ $\mathbf{0}$
NSEI $\mathbf{4.770}$ $\mathbf{0}$ $\mathbf{6.809}$ $\mathbf{0}$ $\mathbf{4.808}$ $\mathbf{0}$ $0.865$ $0.193$ $\mathbf{4.885}$ $\mathbf{0}$
KS11 $\mathbf{3.348}$ $\mathbf{0}$ $\mathbf{3.532}$ $\mathbf{0}$ $\mathbf{2.876}$ $\mathbf{0.002}$ $\mathbf{2.347}$ $\mathbf{0.009}$ $\mathbf{3.363}$ $\mathbf{0}$
BVSP $\mathbf{1.782}$ $\mathbf{0.037}$ $\mathbf{4.378}$ $\mathbf{0}$ $\mathbf{2.217}$ $\mathbf{0.013}$ $\mathbf{2.042}$ $\mathbf{0.021}$ $1.537$ $0.062$
HSI $1.461$ $0.072$ $\mathbf{3.727}$ $\mathbf{0}$ $\mathbf{2.915}$ $\mathbf{0.002}$ $\mathbf{2.840}$ $\mathbf{0.002}$ $1.300$ $0.097$
KSE $\mathbf{2.780}$ $\mathbf{0.003}$ $\mathbf{4.719}$ $\mathbf{0}$ $\mathbf{2.019}$ $\mathbf{0.022}$ $1.171$ $0.121$ $\mathbf{2.683}$ $\mathbf{0.004}$
N225 $\mathbf{4.172}$ $\mathbf{0}$ $\mathbf{3.109}$ $\mathbf{0.001}$ $\mathbf{4.467}$ $\mathbf{0}$ $1.416$ $0.078$ $\mathbf{3.036}$ $\mathbf{0.001}$
SSEC $\mathbf{3.797}$ $\mathbf{0}$ $\mathbf{4.616}$ $\mathbf{0}$ $\mathbf{5.387}$ $\mathbf{0}$ $-0.571$ $0.716$ $\mathbf{4.056}$ $\mathbf{0}$
OSEAX $\mathbf{3.656}$ $\mathbf{0}$ $\mathbf{4.902}$ $\mathbf{0}$ $\mathbf{3.825}$ $\mathbf{0}$ $\mathbf{2.520}$ $\mathbf{0.006}$ $\mathbf{3.274}$ $\mathbf{0.001}$
GSPTSE $\mathbf{4.156}$ $\mathbf{0}$ $\mathbf{4.627}$ $\mathbf{0}$ $\mathbf{5.507}$ $\mathbf{0}$ $-0.084$ $0.533$ $\mathbf{4.453}$ $\mathbf{0}$
@@
According to the DM test results in Tables 7, 8 and 9, in all short-term, mid-term and long-term forecasting tasks, the GSPHAR model has significantly higher RV forecasting accuracy compared to all baseline models, including the HAR, VHAR, HAR-KS-RV and GNNHAR models, for the majority of, if not all, the stock market indices. Similar to the GNNHAR model, the GSPHAR model also applies both linear and nonlinear transformations in merging the asymmetric volatility information from both the real and imaginary domains. This indicates that the volatility spillover effect is directional and has a complex nonlinear mechanism. In addition, compared to the GNNHAR model, the application of the magnet Laplacian in the GSPHAR model allows more efficient and effective graph information propagation. The design of the GSPHAR model not only enables it to easily scale up to larger volatility networks, but also allows it to better capture the asymmetric relationship globally.
The GSPHAR model is also significantly better than the v-GSPHAR model for most stock market indices in different forecasting horizons. This indicates that using a positive $q$ in the magnet Laplacian and a shallow neural network to capture the directional effect through both the real and imaginary domains is effective, emphasizing the nonlinear, directional and asymmetric nature of the volatility spillover effect.
@@
The HAR models utilize a simple average to aggregate past RV values, which implies equal importance of the lagged terms within the mid-term and long-term periods. However, the learnable weights of the convolutional filters of the GSPHAR model reveal that the importance of past RV information decays more dynamically rather than uniformly. According to Figure 4, the decaying pattern of the learned weights is different from the pattern of the prespecified weights in HAR models. The learned weights provide useful information on how the impact diminishes within the short-term, mid-term and long-term periods. Compared to the HAR model weights, the convolution filters put more weight on the $t-1$ RV data, the short-term volatility pattern. This implies that the autoregression in the volatility process is more reactive to recent market conditions. From time lag $2$ to time lag $5$, the learned weights of the GSPHAR model decrease from about $0.38$ to $0.1$ while the weights of HAR models remain constant at $0.2$. This suggests that the HAR model potentially underestimates the importance of the more recent volatility pattern at time lag $2$ and overestimates the importance of the volatility pattern from time lag $3$ to time lag $5$ in the mid-term period. In the long-term period, similar conclusions can be made. The learned weights firstly start above the constant weights from time lag $6$ to time lag $9$, then fluctuate around the constant weights from time lag $10$ to time lag $16$ and finally remain smaller than the constant weights from time lag $17$ to time lag $22$. Although the general pattern of the constant weights of the HAR models is reasonable in that they do put more emphasis on more recent volatility patterns, the stepwise decline between each time period and the stability within each time period overlook the nuanced contributions of different time lags. On the other hand, the learnable weights indicate a more sophisticated understanding of how the influence of past RV values fades with time in the volatility autoregression process and provide a more flexible representation that better captures the temporal relationships inherent in the data.
@@
Compared to the GNNHAR which uses a fixed optimizable hyperparameter $l$ to capture the local volatility interdependencies, the GSPHAR model only relies on the precomputed adjacency matrix to capture the volatility spillover effect. This adjacency matrix can be updated easily as its construction is independent of the model training process. The d-GSPHAR is proposed to incorporate the changing volatility spillover networks by updating the adjacency matrix with correlation coefficient matrices, as described in Equation (36). To evaluate the benefits of the d-GSPHAR in RV forecasting, the index-wise MAE results produced by GSPHAR and d-GSPHAR with different forecast horizons $H$ are presented in Table 10. Although the MAE results between GSPHAR and d-GSPHAR are close in the short-term forecasting task, the d-GSPHAR still delivers better performances in the mid-term and long-term RV forecasting.
@@
BFX $\mathbf{0.168}$ $0.170$ $0.211$ $\mathbf{0.211}$ $0.242$ $\mathbf{0.241}$
@@
SSEC $0.163$ $\mathbf{0.163}$ $0.199$ $\mathbf{0.199}$ $0.243$ $\mathbf{0.242}$
@@ 6 Conclusion
In this paper, three novel multivariate volatility forecasting models, the v-GSPHAR, the GSPHAR and the d-GSPHAR models, are proposed. Key contributions are summarized below:
  • The proposed models are the first attempts to bring the GSP techniques to classical HAR models and achieve significantly better results compared to existing HAR-type baseline models. The proposed models do not handle the volatility spillover effect as a secondary adjustment for overlooking the interconnection between different assets. Rather, the proposed models incorporate the volatility spillover effect as an integral component of the model’s design, which ensures that this phenomenon is embedded to effectively capture the complex interdependencies within financial systems. Based on GSP techniques with the magnet Laplacian, these models extend the traditional HAR framework to capture volatility spillover effects across global financial markets and improve volatility forecasting accuracy. The empirical results demonstrate that the GSPHAR model consistently outperforms baseline models, including the univariate HAR model and its multivariate extensions, and the v-GSPHAR model in short-term, mid-term and long-term multivariate RV forecasting problems.
  • The proposed model provides new insights into the volatility dynamics. The convolution filters disclose the more detailed volatility decaying process. The decaying volatility autoregression does not only exist between the short-term, mid-term and long-term periods, but it also occurs within each historical period. Besides, the d-GSPHAR model novelly integrates the changing volatility spillover effect into the multivariate RV forecasting problem. The dynamic volatility network design allows the proposed model to capture the changing volatility spillover effect as detected in Diebold and Yilmaz (2012); Antonakakis et al. (2018); Gong et al. (2021), which also enhances the forecasting accuracy.
Compared to the existing multivariate HAR model extensions, the GSPHAR model offers a scalable and efficient solution for incorporating the volatility spillover effect without succumbing to common issues, such as insufficient scalability, time-consuming hyperparameter tuning, over-localization and fixed correlations. These findings suggest that the appropriate application of GSP techniques holds great potential for multivariate volatility forecasting and can provide valuable insights for market participants in financial risk management and decision-making. Furthermore, as mentioned in previous sections, how to efficiently and effectively integrate the dynamic volatility spillover effect into the GSPHAR-related forecasting models is an interesting problem for future research.
7 Disclosure of Interest
No conflict of interest is to be declared.
8 Disclosure of Funding
No funding was received.
9 Data Availability Statement
Data were downloaded from Oxford-man Institute’s realized library. The authors confirm that the data supporting the findings of this study are available within the supplementary materials of the paper.

### v2

@@ 4.3 Experiment Results
Index DM Stats p-Value DM Stats p-Value DM Stats p-Value DM Stats p-Value
FCHI $\mathbf{3.056}$ $\mathbf{0.001}$ $\mathbf{6.550}$ $\mathbf{0}$ $\mathbf{4.696}$ $\mathbf{0}$ $\mathbf{2.066}$ $\mathbf{0.019}$
AEX $\mathbf{3.243}$ $\mathbf{0.001}$ $\mathbf{4.406}$ $\mathbf{0}$ $\mathbf{4.004}$ $\mathbf{0}$ $\mathbf{2.302}$ $\mathbf{0.011}$
BFX $\mathbf{2.278}$ $\mathbf{0.011}$ $\mathbf{3.548}$ $\mathbf{0}$ $\mathbf{2.855}$ $\mathbf{0.002}$ $\mathbf{2.890}$ $\mathbf{0.002}$
STOXX50E $\mathbf{3.488}$ $\mathbf{0}$ $\mathbf{6.351}$ $\mathbf{0}$ $\mathbf{4.853}$ $\mathbf{0}$ $\mathbf{2.968}$ $\mathbf{0.002}$
IBEX $\mathbf{2.729}$ $\mathbf{0.003}$ $\mathbf{7.057}$ $\mathbf{0}$ $\mathbf{5.014}$ $\mathbf{0}$ $\mathbf{3.364}$ $\mathbf{0}$
GDAXI $\mathbf{3.603}$ $\mathbf{0}$ $\mathbf{5.551}$ $\mathbf{0}$ $\mathbf{4.983}$ $\mathbf{0}$ $\mathbf{2.795}$ $\mathbf{0.003}$
AORD $\mathbf{2.529}$ $\mathbf{0.006}$ $\mathbf{2.119}$ $\mathbf{0.017}$ $\mathbf{2.913}$ $\mathbf{0.002}$ $\mathbf{2.724}$ $\mathbf{0.003}$
FTSE $\mathbf{3.048}$ $\mathbf{0.001}$ $\mathbf{3.396}$ $\mathbf{0}$ $\mathbf{3.567}$ $\mathbf{0}$ $\mathbf{3.254}$ $\mathbf{0.001}$
MXX $0.610$ $0.271$ $\mathbf{3.835}$ $\mathbf{0}$ $1.117$ $0.132$ $\mathbf{7.176}$ $\mathbf{0}$
IXIC $\mathbf{2.009}$ $\mathbf{0.022}$ $\mathbf{2.575}$ $\mathbf{0.005}$ $\mathbf{2.313}$ $\mathbf{0.010}$ $\mathbf{2.496}$ $\mathbf{0.006}$
SSMI $\mathbf{3.234}$ $\mathbf{0.001}$ $\mathbf{4.484}$ $\mathbf{0}$ $\mathbf{3.569}$ $\mathbf{0}$ $-0.149$ $0.559$
SPX $\mathbf{3.327}$ $\mathbf{0}$ $\mathbf{3.965}$ $\mathbf{0}$ $\mathbf{3.998}$ $\mathbf{0}$ $1.370$ $0.085$
RUT $\mathbf{1.667}$ $\mathbf{0.048}$ $\mathbf{3.054}$ $\mathbf{0.001}$ $\mathbf{2.246}$ $\mathbf{0.012}$ $\mathbf{2.920}$ $\mathbf{0.002}$
DJI $\mathbf{3.438}$ $\mathbf{0}$ $\mathbf{3.570}$ $\mathbf{0}$ $\mathbf{4.055}$ $\mathbf{0}$ $1.586$ $0.056$
BSESN $\mathbf{4.805}$ $\mathbf{0}$ $\mathbf{6.721}$ $\mathbf{0}$ $\mathbf{4.738}$ $\mathbf{0}$ $0.762$ $0.223$
NSEI $\mathbf{4.770}$ $\mathbf{0}$ $\mathbf{6.809}$ $\mathbf{0}$ $\mathbf{4.808}$ $\mathbf{0}$ $0.865$ $0.193$
KS11 $\mathbf{3.348}$ $\mathbf{0}$ $\mathbf{3.532}$ $\mathbf{0}$ $\mathbf{2.876}$ $\mathbf{0.002}$ $\mathbf{2.347}$ $\mathbf{0.009}$
BVSP $\mathbf{1.782}$ $\mathbf{0.037}$ $\mathbf{4.378}$ $\mathbf{0}$ $\mathbf{2.217}$ $\mathbf{0.013}$ $\mathbf{2.042}$ $\mathbf{0.021}$
HSI $1.461$ $0.072$ $\mathbf{3.727}$ $\mathbf{0}$ $\mathbf{2.915}$ $\mathbf{0.002}$ $\mathbf{2.840}$ $\mathbf{0.002}$
KSE $\mathbf{2.780}$ $\mathbf{0.003}$ $\mathbf{4.719}$ $\mathbf{0}$ $\mathbf{2.019}$ $\mathbf{0.022}$ $1.171$ $0.121$
N225 $\mathbf{4.172}$ $\mathbf{0}$ $\mathbf{3.109}$ $\mathbf{0.001}$ $\mathbf{4.467}$ $\mathbf{0}$ $1.416$ $0.078$
SSEC $\mathbf{3.797}$ $\mathbf{0}$ $\mathbf{4.616}$ $\mathbf{0}$ $\mathbf{5.387}$ $\mathbf{0}$ $-0.571$ $0.716$
OSEAX $\mathbf{3.656}$ $\mathbf{0}$ $\mathbf{4.902}$ $\mathbf{0}$ $\mathbf{3.825}$ $\mathbf{0}$ $\mathbf{2.520}$ $\mathbf{0.006}$
GSPTSE $\mathbf{4.156}$ $\mathbf{0}$ $\mathbf{4.627}$ $\mathbf{0}$ $\mathbf{5.507}$ $\mathbf{0}$ $-0.084$ $0.533$
@@
According to the DM test results in Tables 7, 8 and 9, in all short-term, mid-term and long-term forecasting tasks, the GSPHAR model has significantly higher RV forecasting accuracy compared to all baseline models, including the HAR, VHAR, HAR-KS-RV and GNNHAR models, for the majority of, if not all, the stock market indices. The significant superiority of the GSPHAR model demonstrates that, compared to other HAR model extensions, the GSPHAR model better incorporates the directional and nonlinear volatility spillover effect into the HAR framework. It also proves the effectiveness of including the spectral domain analysis in the spatio-temporal forecasting framework. Overall, the GSPHAR model can easily scale up to larger volatility networks, and it better captures the volatility spillover effect at the same time.
@@
The HAR models utilize a simple average to aggregate past RV values, which implies equal importance of the lagged terms within the mid-term and long-term periods. However, the learnable weights of the convolutional filters of the GSPHAR model reveal that the importance of past RV information decays more dynamically and smoothly instead of in a step-wise manner. According to Figure 3, compared to the HAR model weights, the convolution filters put more weight on the $t-1$ RV data, the short-term volatility pattern. This implies that the auto-regression in the volatility process is more reactive to recent market conditions. From time lag $2$ to time lag $5$, the learned weights of the GSPHAR model decrease from about $0.38$ to $0.1$ while the weights of HAR models remain constant at $0.2$. This suggests that the HAR model potentially underestimates the importance of the more recent lag $2$ volatility pattern and overestimates the importance of the volatility pattern of the subsequent lags in the mid-term period. A similar pattern can be detected in the long-term, where the learned weights initially exceed, then fluctuate around, and finally fall below the constant HAR weights. Although the general pattern of the constant weights of the HAR models is reasonable in that they do put more emphasis on more recent volatility patterns, the stepwise decline between each time period and the stability within each time period overlook the nuanced contributions of different time lags. On the other hand, the learnable weights indicate a more sophisticated understanding of how the influence of past RV values fades with time in the volatility auto-regression process and provide a more flexible representation that better captures the temporal relationships inherent in the data.
@@
Compared to the GNNHAR which uses a fixed optimizable hyperparameter $l$ to capture the local volatility interdependencies, the GSPHAR model only relies on the precomputed adjacency matrix to capture the volatility spillover effect. This adjacency matrix can be updated easily as its construction is independent of the model training process. The d-GSPHAR is proposed to incorporate the changing volatility spillover networks by updating the adjacency matrix with correlation coefficient matrices, as described in Equation (17). To evaluate the benefits of the d-GSPHAR in RV forecasting, the index-wise MAE results produced by GSPHAR, the best model in previous comparisons, and d-GSPHAR with different forecasting horizons $h$ are presented in Table 10. Although the MAE results between GSPHAR and d-GSPHAR are close in the short-term forecasting task, the d-GSPHAR improves the RV forecasting performances in the mid-term and long-term RV forecasting. These findings emphasizes the benefits of employing a dynamic graph structure to capture volatility spillover effects and improve the forecasting accuracy.
@@
BFX $\mathbf{0.168}$ $0.170$ $\mathbf{0.211}$ $0.211$ $0.242$ $\mathbf{0.241}$
@@
SSEC $0.163$ $\mathbf{0.163}$ $\mathbf{0.199}$ $0.199$ $0.243$ $\mathbf{0.242}$
@@ 5 Conclusion
This study examines the interconnections of global financial markets by analyzing volatility spillovers. It proposes a novel RV forecasting framework that combines the HAR model and GSP techniques based on the magnetic Laplacian. Results from the empirical tests on data from $24$ global stock market indices show that, compared to baseline models, the GSPHAR model under the proposed GSPHAR framework significantly improves the RV forecasting accuracy. Unlike existing approaches where the volatility spillover effect is regarded as exogenous variables representing additional spatial information in the HAR framework to perform RV forecast, the proposed GSPHAR framework treats the volatility spillover effect as an integral component of the model’s design. The proposed framework uses GFT techniques to incorporate the volatility interrelationships into the HAR model inputs through spatial and spectral domain transformation. In addition, the directionality and nonlinearity of the volatility spillover effect are handled in the spatial domain with a shallow neural network. The superior experiment results indicate that the proposed GSPHAR framework is more effective than the HAR model and its extensions. Furthermore, the GSPHAR framework employs convolution filters with learnable parameters to reveal the more dynamic decaying pattern of the volatility auto-regression process. Besides, the d-GSPHAR model novelly integrates the changing volatility spillover effect into the RV forecasting problem. The dynamic volatility network design allows the proposed model to capture the changing volatility interconnection pattern, which also enhances the forecasting accuracy.
An interesting future research direction can be how to efficiently and effectively design the dynamic volatility spillover graph. The dynamic graph structure should be timely adjusted in response to the evolving relationships among markets or changes in market behavior.
@@ Appendix A RV Time Series Data Statistics
Appendix B Volatility Spillover Effect Matrix Formulation Under DY Framework
Suppose a covariance stationary multi-asset volatility time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consists of $N$ assets. A $p$-order vector autoregressive model (VAR($p$)) for these time series can be represented as:
$\mathbf{v}_{t}=\sum_{i=1}^{\infty}\mathbf{B}_{i}\bm{\epsilon}_{t-i},$ (20)